# Towards Robust and Controllable Text-to-Motion via Masked Autoregressive Diffusion

### v1

@@ Abstract.
Generating 3D human motion from text descriptions remains challenging due to the diverse and complex nature of human motion. While existing methods excel within the training distribution, they often struggle with out-of-distribution motions, limiting their applicability in real-world scenarios. Existing VQVAE-based methods often fail to represent novel motions faithfully using discrete tokens, which hampers their ability to generalize beyond seen data. Meanwhile, diffusion-based methods operating on continuous representations often lack fine-grained control over individual frames. To address these challenges, we propose a robust motion generation framework MoMADiff, which combines masked modeling with diffusion processes to generate motion using frame-level continuous representations. Our model supports flexible user-provided keyframe specification, enabling precise control over both spatial and temporal aspects of motion synthesis. MoMADiff demonstrates strong generalization capability on novel text-to-motion datasets with sparse keyframes as motion prompts. Extensive experiments on two held-out datasets and two standard benchmarks show that our method consistently outperforms state-of-the-art models in motion quality, instruction fidelity, and keyframe adherence.
Human Motion Generation, Masked Modeling, Diffusion Model
@@ 3.2. Training Latent Motion Transformer
VPoser-t (Pavlakos et al., 2019) $75.6$ $48.6$ $9.3$ $1.430$ $8.336$
ACTOR (Petrovich et al., 1 10) $65.3$ $41.0$ $7.0$ $0.341$ $9.569$
MLD (Chen et al., 2023) $\boldsymbol{14.7}$ $8.9$ $5.1$ $0.017$ $9.554$
ParCo (Zou et al., 2024) $53.4$ $38.1$ $7.3$ $0.021{{}^{\pm.000}}$ $9.388{{}^{\pm.078}}$
MotionGPT (Zhang et al., 3 24) $49.7$ $33.2$ $7.7$ $0.089{{}^{\pm.001}}$ $9.653{{}^{\pm.070}}$
MoMask (Guo et al., 2024) $31.3$ $19.2$ $6.3$ $0.020{{}^{\pm.000}}$ $9.616{{}^{\pm.090}}$
Ours $16.4$ $\boldsymbol{3.3}$ $\boldsymbol{3.5}$ $\boldsymbol{0.001}{{}^{\pm.000}}$ $\boldsymbol{9.481}{{}^{\pm.080}}$
@@ 3.3. Inference-time Strategies
ParCo (Zou et al., 2024) $115.3$ $81.4$ $9.2$ $3.932{{}^{\pm.007}}$ $14.311{{}^{\pm.126}}$
MotionGPT (Zhang et al., 3 24) $102.7$ $71.1$ $9.7$ $6.514{{}^{\pm.017}}$ $14.148{{}^{\pm.119}}$
MoMask (Guo et al., 2024) $63.8$ $37.9$ $9.1$ $1.688{{}^{\pm.005}}$ $14.935{{}^{\pm.096}}$
@@
ParCo (Zou et al., 2024) $133.0$ $91.9$ $21.3$ $2.271{{}^{\pm.034}}$ $12.302{{}^{\pm.108}}$
MotionGPT (Zhang et al., 3 24) $163.3$ $114.7$ $22.0$ $9.689{{}^{\pm.103}}$ $10.132{{}^{\pm.087}}$
MoMask (Guo et al., 2024) $99.3$ $58.9$ $21.3$ $1.442{{}^{\pm.024}}$ $12.797{{}^{\pm.099}}$
@@ 4.1.1. Datasets
HumanML3D dataset collects the motions contains 14,616 human motions from AMASS (Mahmood et al., 2019) dataset and HumanAct12 (Guo et al., 2020), with 44,970 textual descriptions. KIT-ML dataset contains 3,911 motions from KIT (Mandery et al., 2015) and CMU (car, [n. d.]), and includes 6,278 descriptions. IDEA400 dataset is the largest subset apart from the AMASS dataset, which contains 12,042 motion sequences with one text description each. Kungfu includes 1,032 Chinese kungfu motion sequences with one semantic text description per sequence. More details can be found in the supplementary.
@@ 4.1.2. Evaluation Metrics
We adopt motion-text evaluator from (Guo et al., 2022a) and use the following evaluation metrics. (1) R-Precision measures text-motion alignment by computing Euclidean distances between a motion feature and 32 candidate text features. We report top-1, top-2, and top-3 retrieval accuracies. (2) Frechet Inception Distance (FID) (Heusel et al., 2017) evaluates the distributional similarity between generated and real motions, based on features extracted by the motion encoder. (3) Multimodal Distance (MMD) calculates the average Euclidean distance between motion features and their corresponding text features. (4) Diversity (DIV) assesses the variety in generated motions by computing the average Euclidean distance between 300 randomly sampled pairs.
@@ 4.1.3. Implementation Details
Masked Autoregressive Diffusion Model. The transformer consists of 16 layers, each with 8 attention heads. The hidden dimension is 1024, and the output condition token dimension is 512. The diffusion head is implemented as a 4-layer MLP. We use a learning rate of 0.0001 with a linear warm-up over the first 2000 iterations. For HumanML3D, the model is trained for 600 epochs, with a decay factor of 0.1 applied at epoch 400. For the KIT-ML dataset, we train for 1500 epochs with decay at epoch 1200. To stabilize the training process, we apply an exponential moving average (EMA) to model parameters with a decay rate of 0.999. For the DDPM variant, we use 50 diffusion steps. For the DDIM variant, we train with 1000 diffusion steps and use 100 steps during inference.

### v2

@@ Abstract.
Generating 3D human motion from text descriptions remains challenging due to the diverse and complex nature of human motion. While existing methods excel within the training distribution, they often struggle with out-of-distribution motions, limiting their applicability in real-world scenarios. Existing VQVAE-based methods often fail to represent novel motions faithfully using discrete tokens, which hampers their ability to generalize beyond seen data. Meanwhile, diffusion-based methods operating on continuous representations often lack fine-grained control over individual frames. To address these challenges, we propose a robust motion generation framework MoMADiff, which combines masked modeling with diffusion processes to generate motion using frame-level continuous representations. Our model supports flexible user-provided keyframe specification, enabling precise control over both spatial and temporal aspects of motion synthesis. MoMADiff demonstrates strong generalization capability on novel text-to-motion datasets with sparse keyframes as motion prompts. Extensive experiments on two held-out datasets and two standard benchmarks show that our method consistently outperforms state-of-the-art models in motion quality, instruction fidelity, and keyframe adherence. The code is available at: https://github.com/zzysteve/MoMADiff
Human Motion Generation, Text-to-Motion, Masked Modeling, Diffusion Model
© Owner/Author — ACM 2025. This is the author’s version of the work. It is posted here for your personal use. Not for redistribution. The definitive Version of Record was published in Proceedings of the 33rd ACM International Conference on Multimedia, http://dx.doi.org/10.1145/3746027.3754748.
conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland
@@ 3.2. Training Latent Motion Transformer
VPoser-t (Pavlakos et al., 2019) 75.675.675.6 48.648.648.6 9.39.39.3 1.4301.4301.430 8.3368.3368.336
ACTOR (Petrovich et al., 1 10) 65.365.365.3 41.041.041.0 7.07.07.0 0.3410.3410.341 9.5699.5699.569
MLD (Chen et al., 2023) $\boldsymbol{14.7}$ 8.98.98.9 5.15.15.1 0.0170.0170.017 9.5549.5549.554
ParCo (Zou et al., 2024) 53.453.453.4 38.138.138.1 7.37.37.3 $0.021{{}^{\pm.000}}$ $9.388{{}^{\pm.078}}$
MotionGPT (Zhang et al., 3 24) 49.749.749.7 33.233.233.2 7.77.77.7 $0.089{{}^{\pm.001}}$ $9.653{{}^{\pm.070}}$
MoMask (Guo et al., 2024) 31.331.331.3 19.219.219.2 6.36.36.3 $0.020{{}^{\pm.000}}$ $9.616{{}^{\pm.090}}$
Ours 16.416.416.4 $\boldsymbol{3.3}$ $\boldsymbol{3.5}$ $\boldsymbol{0.001}{{}^{\pm.000}}$ $\boldsymbol{9.481}{{}^{\pm.080}}$
@@ 3.3. Inference-time Strategies
ParCo (Zou et al., 2024) 115.3115.3115.3 81.481.481.4 9.29.29.2 $3.932{{}^{\pm.007}}$ $14.311{{}^{\pm.126}}$
MotionGPT (Zhang et al., 3 24) 102.7102.7102.7 71.171.171.1 9.79.79.7 $6.514{{}^{\pm.017}}$ $14.148{{}^{\pm.119}}$
MoMask (Guo et al., 2024) 63.863.863.8 37.937.937.9 9.19.19.1 $1.688{{}^{\pm.005}}$ $14.935{{}^{\pm.096}}$
@@
ParCo (Zou et al., 2024) 133.0133.0133.0 91.991.991.9 21.321.321.3 $2.271{{}^{\pm.034}}$ $12.302{{}^{\pm.108}}$
MotionGPT (Zhang et al., 3 24) 163.3163.3163.3 114.7114.7114.7 22.022.022.0 $9.689{{}^{\pm.103}}$ $10.132{{}^{\pm.087}}$
MoMask (Guo et al., 2024) 99.399.399.3 58.958.958.9 21.321.321.3 $1.442{{}^{\pm.024}}$ $12.797{{}^{\pm.099}}$
@@ 4.1.1. Datasets
HumanML3D dataset collects the motions contains 14,616 human motions from AMASS (Mahmood et al., 2019) dataset and HumanAct12 (Guo et al., 2020), with 44,970 textual descriptions. KIT-ML dataset contains 3,911 motions from KIT (Mandery et al., 2015) and CMU (CMU Graphics Lab, [n. d.]), and includes 6,278 descriptions. IDEA400 dataset is the largest subset apart from the AMASS dataset, which contains 12,042 motion sequences with one text description each. Kungfu includes 1,032 Chinese kungfu motion sequences with one semantic text description per sequence. More details can be found in the supplementary.
@@ 4.1.2. Evaluation Metrics
We adopt motion-text evaluator from (Guo et al., 2022a) and use the following evaluation metrics. (1) R-Precision measures text-motion alignment by computing Euclidean distances between a motion feature and 32 candidate text features. We report top-1, top-2, and top-3 retrieval accuracies. (2) Frechet Inception Distance (FID) (Heusel et al., 2017) evaluates the distributional similarity between generated and real motions, based on features extracted by the motion encoder. (3) Multimodal Distance (MM-Dist) calculates the average Euclidean distance between motion features and their corresponding text features. (4) Diversity (DIV) assesses the variety in generated motions by computing the average Euclidean distance between 300 randomly sampled pairs.
@@ 4.1.3. Implementation Details
Masked Autoregressive Diffusion Model. The transformer consists of 16 layers, each with 8 attention heads. The hidden dimension is 1024, and the output condition token dimension is 512. The diffusion head is implemented as a 4-layer MLP. We use a learning rate of 0.0001 with a linear warm-up over the first 2000 iterations. For HumanML3D, the model is trained for 600 epochs, with a decay factor of 0.1 applied at epoch 400. For the KIT-ML dataset, we train for 1500 epochs with decay at epoch 1200. To stabilize the training process, we apply an exponential moving average (EMA) to model parameters with a decay rate of 0.999. For the DDPM (Ho et al., 2020) variant, we use 50 diffusion steps. For the DDIM variant, we train with 1000 diffusion steps and use 100 steps during inference.
@@ 5. Conclusion
Appendix
Supplementary Material
Appendix A More Ablation Studies
A.1. Key Components
The components of MoMADiff are intentionally designed to be closely integrated. To evaluate the contribution of each component, we performed an ablation study by selectively removing or replacing them. Specifically, we trained the following three variants:
Without VAE: We removed the VAE component, which is responsible for modeling the spatial structure and local motion priors. In this setting, the model directly predicts the motion sequence $X$ in the observation space, instead of modeling its latent representation.
Without Diffusion Head: We removed the diffusion head, so the Transformer directly models the latent variable $Z$ without first producing condition vectors for diffusion.
Transformer Only: We removed both the VAE and diffusion head, leaving only a Transformer. In this case, the model takes masked motion sequences as input and recursively predicts the motion $X$ in an autoregressive manner.
The results are presented in Table I. We observe that the inclusion of the VAE provides useful motion priors, leading to improved performance compared to the baseline without the VAE and diffusion head. However, it remains challenging for the Transformer to predict directly in the continuous motion space. By introducing the diffusion head to handle the generation of continuous representations, the Transformer can instead focus on modeling temporal dependencies. The diffusion process produces more accurate and reliable motion representations, leading to better generation results.
A.2. Inference Modes
We use a cosine scheduler to determine the number of frames to be predicted at each iteration, selecting them randomly from the motion sequence. Additionally, we explore several alternative inference modes to better understand how different generation orders affect performance, as illustrated in Figure I.
Keyframe Mode. In this mode, the masking ratio at the $i$-th step is determined by the function $y=cos(\frac{\pi}{2}\cdot\frac{i}{R})$, where $R$ is the total number of steps. The motion embeddings to be predicted are randomly selected from the sequence, simulating a sparse keyframe first generation process.
Linear Mode. This mode uses a linear function $y=1-\frac{i}{R}$ to control the number of masked frames at each step. The motion embeddings to be predicted are selected sequentially from the beginning to the end of the sequence, following a next-set generation strategy.
Bidirectional Linear. Similar to the Linear mode, this approach also uses $y=1-\frac{i}{R}$ to schedule the masking ratio. However, instead of predicting frames in a single direction, the motion embeddings are selected symmetrically from both the beginning and end of the sequence. Generation progresses inward from both sides in a bidirectional next-set manner.
Quantitative results are presented in Table III. As shown, keyframe mode consistently demonstrates superior performance compared to the other inference strategies.
A.3. Depth of the Diffusion Head
We conduct ablation studies on the number of layers in the diffusion head, with results summarized in Table IV. All experiments are based on the DDPM model using 50 diffusion steps. We observe that the best performance is achieved with four layers of diffusion blocks. Using fewer blocks may limit the model’s capacity to capture motion dynamics, while increasing the number of layers can lead to gradient vanishing issues, likely due to our use of a simple MLP-based design without skip connections.
Additionally, we experimented with the diffusion head architecture proposed in (Li et al., 2024), originally designed for image generation. However, its skip connection structure did not yield satisfactory results in our motion generation task, suggesting that architectural designs optimized for image domains may not transfer well to motion modeling.
Figure I. Illustration of different types of inference modes.
[…]
Appendix B Inference Speed
Table I. Ablation study on the impact of VAE and Diffusion modules
[…]
0.523$\pm$.003 0.713$\pm$.003 0.807$\pm$.002 0.073$\pm$.004 2.917$\pm$.010 9.711$\pm$.070
0.397$\pm$.003 0.571$\pm$.003 0.673$\pm$.002 1.171$\pm$.014 3.864$\pm$.010 9.181$\pm$.082
0.440$\pm$.003 0.631$\pm$.003 0.736$\pm$.002 0.949$\pm$.017 3.469$\pm$.008 9.616$\pm$.079
0.390$\pm$.003 0.570$\pm$.003 0.679$\pm$.003 1.483$\pm$.014 3.789$\pm$.012 9.511$\pm$.087
MoMADiff supports flexible trade-offs between efficiency and quality by adjusting two factors: the number of recursive steps (R) and the number of DDIM inference steps (I. Steps). To quantify computational cost, we report the Average Inference Time per Sentence (AITS) in Table II.
As shown, increasing either R or I. Steps improve generation quality (e.g., lower FID) but also increase inference time, providing users with the flexibility to balance performance and speed according to practical needs. Under a fast inference setting (I. Steps = 10, R=3), our model already achieves slightly better FID scores than previous baselines with comparable inference time. Furthermore, with more DDIM and recursive steps (e.g., I. Steps = 10, R = 9), our model continues to improve in quality while maintaining competitive efficiency. This design also supports fine-grained control over the generation process, allowing dynamic adjustments between speed and quality.
Table II. Comparison with prior methods on motion generation
[…]
TEMOS 3.734 0.017
T2M 1.067 0.038
MotionDiffuse 0.630 14.740
MDM 0.544 24.740
MLD 0.473 0.217
MotionLCM 0.467 0.030
Ours (I. Steps 10, R=3) 0.329 0.074
Ours (I. Steps 50, R=9) 0.107 0.778
Ours (I. Steps 100, R=9) 0.073 1.483
Table III. Ablation on Inference Mode.
[…]
Top1 Top2 Top3
Keyframe 0.523±.003 0.713±.003 0.807±.002 0.073±.004 2.917±.010 9.711±.070
Linear 0.514±.004 0.707±.003 0.804±.002 0.131±.003 2.938±.009 9.555±.067
Bi-directional Linear 0.512±.003 0.703±.003 0.800±.002 0.108±.005 2.960±.009 9.580±.069
Figure II. Additional motion reconstruction results on out-of-distribution motions using different encoders.
[…]
Table IV. Ablation on diffusion head design
[…]
Top1 Top2 Top3
Ours with 3 layers 0.497±.003 0.689±.002 0.785±.002 0.266±.008 3.148±.011 9.819±.073
Ours with 4 layers 0.522±.003 0.716±.003 0.810±.002 0.134±.004 2.910±.010 9.730±.064
Ours with 6 layers 0.516±.003 0.708±.002 0.802±.002 0.191±.005 2.945±.008 9.737±.079
Diffusion head in (Li et al., 2024) 0.450±.003 0.633±.004 0.735±.003 0.675±.017 3.361±.010 9.114±.070
Appendix C Additional Qualitative Results
C.1. Ours on HumanML3D
We provide qualitative results of our method on HumanML3D dataset in Figure III. Please also visit our project page for more qualitative video results, which include comparisons of motion reconstruction, keyframe-guided generation, and text-to-motion generation on the HumanML3D dataset.
C.2. Reconstruction Results
We further illustrate the reconstruction capability of our continuous motion autoencoder by comparing it against existing VQ-VAE architectures with two samples on the Kungfu dataset. As shown in Figure II, our method achieves closer reconstructions to the ground truth compared to previous encoder-based approaches, demonstrating its superior ability to preserve details of out-of-distribution motions.
Figure III. Qualitative results of text-to-motion generation on the HumanML3D dataset.
[…]
Appendix D More Implementation Details
D.1. Evaluation Metrics
We adopt the motion-text evaluator from (Guo et al., 2022a) and use the following evaluation metrics.
(1) R-Precision Measures text-motion alignment by computing Euclidean distances between a motion feature and 32 candidate text features. We report top-1, top-2, and top-3 retrieval accuracies.
(2) Frechet Inception Distance (FID) (Heusel et al., 2017): Evaluates the distributional similarity between generated and real motions, based on features extracted by the motion encoder.
(3) Multimodal Distance (MMD): Calculates the average Euclidean distance between motion features and their corresponding text features.
(4) Diversity: Assesses the variety in generated motions by computing the average Euclidean distance between 300 randomly sampled pairs.
D.2. Datasets
HumanML3D dataset collects the motions from AMASS (Mahmood et al., 2019) dataset and HumanAct12 (Guo et al., 2020) dataset, which contains 14,616 human motions. The dataset provides 44,970 textual descriptions in total for these motions, with three descriptions for each motion sequence. HumanML3D contains diverse actions including daily activities, sports, acrobatics, and artistry.
KIT-ML dataset contains 3,911 motions from KIT (Mandery et al., 2015) and CMU (CMU Graphics Lab, [n. d.]), and includes 6,278 descriptions.
IDEA400 dataset is the largest subset apart from AMASS dataset, which contains 12,042 motion sequences with one text description each. It contains 400 actions with human self-contact motions and human-object contact motions during walking, standing, and sitting, which examine the detailed modeling capability of the model.
Kungfu includes 1,032 motion sequences with one semantic text description per sequence. It represents a highly challenging out-of-distribution evaluation scenario due to its complex and stylized motion patterns.
IDEA400 and Kungfu are derived from the high-quality, whole-body, large-scale human motion dataset Motion-X. To ensure compatibility with HumanML3D, we extract body-only poses and semantic text descriptions and format the data to match the HumanML3D specification.
Since motions in the HumanML3D dataset have a maximum length of 196 frames, we follow this constraint when evaluating on the IDEA400 and Kungfu datasets by selecting only motion sequences shorter than 196 frames.
  Hardware Environment
  CPU: Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz
  Memory: 256 GB
  GPU: GeForce RTX 3090
  Software Environment
  sys.platform: linux
  Python: 3.10.14 [GCC 11.2.0]
  CUDA available: True
  GPU 0,1: NVIDIA GeForce RTX 3090
  GCC: gcc (Ubuntu 7.5.0-3ubuntu1~18.04) 7.5.0
  PyTorch: 2.4.0+cu121
  PyTorch compiling details: PyTorch built with:
  - GCC 9.3
  - C++ Version: 201703
  - Intel(R) oneAPI Math Kernel Library Version 2022.2-Product Build 20220804 for Intel(R) 64
  architecture applications
  - Intel(R) MKL-DNN v3.4.2
  - OpenMP 201511 (a.k.a. OpenMP 4.5)
  - LAPACK is enabled (usually provided by MKL)
  - NNPACK is enabled
  - CPU capability usage: AVX512
  - CUDA Runtime 12.1
  - NVCC architecture flags: -gencode;arch=compute_50,code=sm_50;-gencode;arch=compute_60,
  code=sm_60;-gencode;arch=compute_70,code=sm_70;-gencode;arch=compute_75,code=sm_75;-gencode;
  arch=compute_80,code=sm_80;-gencode;arch=compute_86,code=sm_86;-gencode;arch=compute_90,
  code=sm_90
  - CuDNN 90.1  (built against CUDA 12.4)
  - Magma 2.6.1
  TorchVision: 0.19.0+cu121
Figure IV. Hardware and software environments.
D.3. Hardware and Software Environments
The hardware and software environments used in our experiments are illustrated in Figure IV. All processes, including training and inference, are conducted on machines with these configurations.